Question 4: (Excel: enter data manually)

How is the hatching of water python eggs influenced by the temperature of the snake’s nest? Researchers assigned newly laid eggs to one of three temperatures: hot, neutral, or cold. Hot duplicates the warmth provided by the mother python. Neutral and cold are cooler, as when the mother is absent. Here are the data on the number of eggs hatched.

Cold       Neutral                Warm

Number of eggs               27          56                         104

Number hatched             16          38                         75

a) Notice that this is not a 2-way table. Summarize the data above in a two-way table where the rows are the number of eggs hatched and the number not hatched.

b) Find the row distributions. Summarize them in a table.

c) Find the column distributions. Summarize them in a table.

d) Make a stacked bar chart that makes the most sense to you. In setting up your chart make sure you choose the data series in rows/columns correctly. Levels of the explanatory variable must be listed along the horizontal axis. You may have to change the rows/columns button until you have it right.
Answer #1

(a)

Not hatched

Hatched

Cold

27-16 = 11

16

Neutral

56-38 = 18

38

Hot

104-75 = 29

75

Mately Not Match Tstals Cold Neltoal t87 24 WArm Total 58 L81 129yloo +7 58X00 31-022

(d)

Lets make conditional distribution of temperature for hatched eggs and then make a bar graph to display this distribution .

16/129=12.40%,

38/129=29.46%,

75/129=58.14%,

80 58.14% 60 29.46 % 40 12.40% 20 0 cold neutral hot temperature for hatched eggs percent